[clas_members] [New Logentry] RC Daily Summary - Sunday

daria at jlab.org daria at jlab.org
Sun Sep 11 16:00:02 EDT 2022

Logentry Text:
Dear all, 

A slightly more eventful past 24h.

Some DAQ problems during the day yesterday and a proper tantrum this morning -- needed a full ROC reboot by Sergey to fix. 

Yesterday evening the NMR started showing a slight decrease in polarisation amplitude. I tried to adjust but the manual step motor control did not want to be touched -- frequency went wild for a while, finally managed to calm it down to 140.49 GHz, a notch up from previous 140.48 GHz. Seems optimum for the target condition (just a bit above 30%, negative pol), but NMR still shows a gradually falling polarisation. Decided to run overnight with this. https://logbooks.jlab.org/entry/4039990 https://logbooks.jlab.org/entry/4040022

Harut has been analysing the data hot-off-the-press, that Mohammed's been cooking (thank you both!), and observed a significant drop in the polarisation. It was decided this morning to anneal the target. Maxwell, Victoria, Tsuneo, Pushpa did the anneal at 1pm today, successfully completed -- the target is now polarising. We will start with negative polarisation.

Just before 9pm last night the RICH Sector 1 hardware interlocks packed in: https://logbooks.jlab.org/entry/4040015
Valery tried to restart them remotely but total communication failure. Emailed Detector Support Group, but 'twas a tad late on a Saturday night to expect an immediate response. Valery came in and did a rapid-access power-cycle on the crate downstairs in the hall (thank you muchly!): that fixed the problem. RICH no longer at risk of catching fire. https://logbooks.jlab.org/entry/4040028
This morning a RICH glitch (but was fixed with the RICH recovery procedure).

MCC took about 30min to bring the beam back after the Controlled Access (magnets ramping up) -- weird, just now it took 3min.

We flipped the half-wave plate to OUT at 10.30pm last night.  

During the night there were a few beam-down incidents (lost about an hour) and a few trips, also some drifting beam positions which the shifters caught.

Nathan added a correction for the Faraday Cup beam-charge asymmetry to the helicity asymmetry: https://logbooks.jlab.org/entry/4039987
There are changes in the asymmetries reflecting the HW-plate changes and the target pol reversal, but the results are still looking a bit odd. We're trying to make sense of them.

See you at the RC meeting on Monday at 2pm! 

Have a good rest of Sunday, 

Your RC

******** Stats since the start of ND3 with ELMO ********
Half-wave plate OUT, target polarisation POSITIVE:

16902 - 16908: details in the previous RC summary.

TOTAL triggers from "good" runs: 467 M

Half-wave plate IN, target polarisation POSITIVE:

16910 - 16922: details in the previous RC summary.

TOTAL triggers from "good" runs: 498 M (slightly lower target pol than the set with OUT half-wave plate).

Half-wave plate IN, target polarisation NEGATIVE:

16923: 105.3 M (target pol increasing during this run).
16924: 100.0 M (pinned)
16925: 44.0 M (Brief RICH glitch, DAQ crash at the end)
16927: 100.0 M
16928: 100.0 M
16929: 27.8 M (pinned, DAQ crash at end)
16931: 81.6 M (40min of no beam at end, about an hour of fiddling with target pol in middle)
16832: 79.4 M 
TOTAL triggers from "good" runs: 638 M

Half-wave plate OUT, target polarisation NEGATIVE:

16933: 103.9 M (pinned)
16934: 100.0 M (beam position drifts towards end)
16935: 100.4 M
16944: 30.8 M

Stopped to anneal target -- polarisation clearly dropped significantly in the past day or so.

TOTAL triggers from "good" runs: 335 M


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4040235
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220911/1422be50/attachment.html>

More information about the clas_members mailing list